OpenStudy (anonymous):

find dy/dx by implicit differentiation: xy+y^2=4

OpenStudy (bahrom7893):

x*y'+y+2yy'=0 xy'+2yy'=-y y'(x+2y)=-y y' = -y/(x+2y)

OpenStudy (bahrom7893):

dy/dx = -y/(x+2y)